# Efficient Sparse Coding with the Adaptive Locally Competitive Algorithm for Speech Classification

### v1

@@ Abstract
Researchers are exploring novel computational paradigms such as sparse coding and neuromorphic computing to bridge the efficiency gap between the human brain and conventional computers in complex tasks. A key area of focus is neuromorphic audio processing. While the Locally Competitive Algorithm has emerged as a promising solution for sparse coding, offering potential for real-time and low-power processing on neuromorphic hardware, its applications in neuromorphic speech classification have not been thoroughly studied. The Adaptive Locally Competitive Algorithm builds upon the Locally Competitive Algorithm by dynamically adjusting the modulation parameters of the filter bank to fine-tune the filters’ sensitivity. This adaptability enhances lateral inhibition, improving reconstruction quality, sparsity, and convergence time, which is crucial for real-time applications. This paper demonstrates the potential of the Locally Competitive Algorithm and its adaptive variant as robust feature extractors for neuromorphic speech classification. Results show that the Locally Competitive Algorithm achieves better speech classification accuracy at the expense of higher power consumption compared to the LAUSCHER cochlea model used for benchmarking. On the other hand, the Adaptive Locally Competitive Algorithm mitigates this power consumption issue without compromising the accuracy. The dynamic power consumption is reduced to a range of 0.004 to 13 milliwatts on neuromorphic hardware, three orders of magnitude less than setups using Graphics Processing Units. These findings position the Adaptive Locally Competitive Algorithm as a compelling solution for efficient speech classification systems, promising substantial advancements in balancing speech classification accuracy and power efficiency.
@@ II-A2 Locally Competitive Algorithm (LCA) for Audio
where $\tau$ is the time constant of each neuron, $\boldsymbol{p}$ is the input signal projection on the dictionary, i.e., $\boldsymbol{p}=\boldsymbol{D}^{T}\boldsymbol{s}$, $\boldsymbol{v}$ is the membrane potential vector, and $\boldsymbol{I}$ is the identity matrix. Basically, the evolution of $\boldsymbol{v}$ over time depends on the input intensity $\boldsymbol{p}$ and on $-\boldsymbol{v}$ which makes these neurons behave like leaky integrators. Membrane potentials exceeding the threshold $\lambda$ –––the same as the Lagrange multiplier in (2)––– produce activations corresponding to the coefficients $\boldsymbol{a}$. Each activated neuron inhibits all others through horizontal connections, where $\boldsymbol{D}^{T}\boldsymbol{D}-\boldsymbol{I}$ are the weights of these connections. For each neuron $m$, the activation $a_{m}$ is a non-linearity $T_{\lambda}$ that can be sigmoidal or –––as we used in this work––– the hard thresholding function applied to the potential $v_{m}$ of neuron $m$:
@@
[…]
@@
$\boldsymbol{D}^{T}=\begin{bmatrix}\boldsymbol{\phi}_{0}[0]&...&\boldsymbol{\phi}_{0}[F_{l}]&\boldsymbol{0_{1\times r}}&0&...\\ \boldsymbol{0_{1\times r}}&\boldsymbol{\phi}_{0}[0]&...&\boldsymbol{\phi}_{0}[F_{l}]&\boldsymbol{0_{1\times r}}&...\\ \boldsymbol{0_{1\times r}}&\boldsymbol{0_{1\times r}}&\boldsymbol{\phi}_{0}[0]&...&\boldsymbol{\phi}_{0}[F_{l}]&...\\ \vdots&&&&&\vdots\\ \boldsymbol{\phi}_{i}[0]&...&\boldsymbol{\phi}_{i}[F_{l}]&\boldsymbol{0_{1\times r}}&0&...\\ \boldsymbol{0_{1\times r}}&\boldsymbol{\phi}_{i}[0]&...&\boldsymbol{\phi}_{i}[F_{l}]&\boldsymbol{0_{1\times r}}&...\\ \boldsymbol{0_{1\times r}}&\boldsymbol{0_{1\times r}}&\boldsymbol{\phi}_{i}[0]&...&\boldsymbol{\phi}_{i}[F_{l}]&...\\ \vdots&&&&&\vdots\\ \boldsymbol{\phi}_{k}[0]&...&\boldsymbol{\phi}_{k}[F_{l}]&\boldsymbol{0_{1\times r}}&0&...\\ \boldsymbol{0_{1\times r}}&\boldsymbol{\phi}_{k}[0]&...&\boldsymbol{\phi}_{k}[F_{l}]&\boldsymbol{0_{1\times r}}&...\\ \boldsymbol{0_{1\times r}}&\boldsymbol{0_{1\times r}}&\boldsymbol{\phi}_{k}[0]&...&\boldsymbol{\phi}_{k}[F_{l}]&...\\ \end{bmatrix},$ (9)
@@ II-A3 Adaptive LCA (ALCA)
where |||| is the ”OR” operator and $\boldsymbol{c}$, $\boldsymbol{b}$, and $\boldsymbol{l}$ are vectors containing the parameters of the filters $c_{i}$, $b_{i}$, and $l_{i}$, respectively. This relationship enabled the implementation of the gradient descent, allowing the gradient flow reaching the membrane potential to backpropagate further towards the Gammachirp’s parameters [23].
@@
$E=\frac{1}{2}||\boldsymbol{\hat{s}}-\boldsymbol{s}||^{2}+\alpha\lambda S(\boldsymbol{a}).$ (12)
@@ II-B Audio Datasets
[…]
@@
We also encode both the HD and SC datasets with LCA, selecting $k=700$ filters to match the number of channels in SHD and SSC. This allows us to compare, in Sec. IV, our work with state-of-the-art methods that use SHD/SSC with all the channels. We choose the values for the filter length $F_{l}$ and the stride $r$ that ensure to have similar analysis window (10 ms) as temporal histograms of SHD and SSC. We set the threshold $\lambda=0.00045$ for HD and $\lambda=0.0007$ for SC to maintain similar sparsity as SHD and SSC. The values used for the dictionary hyperparameters are in Tab. I. The representations of HD and SC with this algorithm are called LCA-HD and LCA-SC, respectively. From these LCA representations we derive ALCA representation by adapting the Gammachirps therefore creating two additional representations ALCA-HD and ALCA-SC. All representations are normalized to a $0-1$ range, ensuring all values fall between 00 and $1$.
@@ III-A Static Features
Figure 3: Best accuracy achieved by fully connected ANNs with both HD and SC datasets with three representations: LAUSCHER S[HD||||SC], LCA -[HD||||SC], and ALCA-[HD||||SC].
@@ III-B Importance of Temporal Features
where #$\#_{\text{c}}(.)$ is the average number of non-zero coefficients in the considered representation (ALCA or LCA) computed from the test set of the considered dataset, and #$\#_{\text{s}}(.)$ is the average spike count of the LAUSCHER representation computed from the test set of the considered dataset.
@@ III-D Spiking Activity
Quantized Model Test Accuracy (%) $83.77\pm 1.2$ $\boldsymbol{94.63\pm 0.8}$ $94.38\pm 1.0$ $68.53\pm 0.4$ $84.44\pm 0.5$ $\boldsymbol{86.40\pm 0.4}$
@@ III-E Power/Energy Efficiency on Loihi 2
The inference on the GPU utilizes (Lava-DL SLAYER), and the its efficiency is measured using nvidia-smi on Linux with kernel version 5.4.0-7634, Python 3.8.1 and PyTorch 2.0.0. Loihi’s efficiency is measured using the power probes of Lava 0.9.0 on Oheogultch board ncl-ext-og-05.
@@
LAUSCHER [41] Feed-forward SNN with adaptive axonal delays $92.45$
LAUSCHER [42] Feed-forward SNN with spatio-temporal filters $92.36$
@@
LAUSCHER [43] RSNN with temporal attention $91.08$
LAUSCHER [44] Feed-forward SNN with random dendritic delays $90.88$

### v4

@@ Abstract
Researchers are exploring novel computational paradigms such as sparse coding and neuromorphic computing to bridge the efficiency gap between the human brain and conventional computers in complex tasks. A key area of focus is neuromorphic audio processing. While the Locally Competitive Algorithm has emerged as a promising solution for sparse coding, offering potential for real-time and low-power processing on neuromorphic hardware, its applications in neuromorphic speech classification have not been thoroughly studied. The Adaptive Locally Competitive Algorithm builds upon the Locally Competitive Algorithm by dynamically adjusting the modulation parameters of the filter bank to fine-tune the filters’ sensitivity. This adaptability enhances lateral inhibition, improving reconstruction quality, sparsity, and convergence time, which is crucial for real-time applications. This paper demonstrates the potential of the Locally Competitive Algorithm and its adaptive variant as robust feature extractors for neuromorphic speech classification. Results show that the Locally Competitive Algorithm achieves better speech classification accuracy at the expense of higher power consumption compared to the LAUSCHER cochlea model used for benchmarking. On the other hand, the Adaptive Locally Competitive Algorithm mitigates this power consumption issue without compromising the accuracy. The dynamic power consumption is reduced to a range of $4$ to $13$ milliwatts on neuromorphic hardware, three orders of magnitude less than setups using Graphics Processing Units. These findings position the Adaptive Locally Competitive Algorithm as a compelling solution for efficient speech classification systems, promising substantial advancements in balancing speech classification accuracy and power efficiency.
@@ II-A2 Locally Competitive Algorithm (LCA) for Audio
where $\tau$ is the time constant of each neuron, $\boldsymbol{p}$ is the input signal projection on the dictionary, i.e., $\boldsymbol{p}=\boldsymbol{D}^{T}\boldsymbol{s}$, $\boldsymbol{v}$ is the membrane potential vector, and $\boldsymbol{I}$ is the identity matrix. Basically, the evolution of $\boldsymbol{v}$ over time depends on the input intensity $\boldsymbol{p}$ and on $\boldsymbol{v}$ which makes these neurons behave like leaky integrators. Membrane potentials exceeding the threshold $\lambda$ –––the same as the Lagrange multiplier in (2)––– produce activations corresponding to the coefficients $\boldsymbol{a}$. Each activated neuron inhibits all others through horizontal connections, where $\boldsymbol{D}^{T}\boldsymbol{D}-\boldsymbol{I}$ are the weights of these connections. For each neuron $m$, the activation $a_{m}$ is a non-linearity $T_{\lambda}$ that can be sigmoidal or –––as we used in this work––– the hard thresholding function applied to the potential $v_{m}$ of neuron $m$:
@@
$\displaystyle\begin{aligned} \boldsymbol{v}[n]=&\frac{\Delta t}{\tau}(\boldsymbol{D}^{T}\boldsymbol{s}-(\boldsymbol{D}^{T}\boldsymbol{D}-\boldsymbol{I})\boldsymbol{a}[n-1])\\ &+(1-\frac{\Delta t}{\tau})\boldsymbol{v}[n-1]\end{aligned}.$ (6)
@@
[…]
@@ II-A3 Adaptive LCA (ALCA)
where $|$ is the ”OR” operator and $\boldsymbol{c}$, $\boldsymbol{b}$, and $\boldsymbol{l}$ are vectors containing the parameters of the filters $c_{i}$, $b_{i}$, and $l_{i}$, respectively. This relationship enabled the implementation of the gradient descent, allowing the gradient flow reaching the membrane potential to backpropagate further towards the Gammachirp’s parameters [23].
@@
[…]
@@ II-B Audio Datasets
Two speech classification tasks are studied. First, the Heidelberg digits dataset (HD) [24] that consists of approximately $10\thinspace{}420$ recordings of spoken digits from zero to nine in both English and German languages including $12$ speakers in total. The sampling rate is $48$ kHz. The dataset is separated into training and test sets. Two speakers are exclusively held for the test set. The remainder of the test set is filled with samples ($5\%$ of the recordings) from speakers also present in the training set. This division allows one to assess a trained network’s ability to generalize across speakers.
The second task is The Google Speech commands dataset (SC), comprising $1$ second long .wav-files with a $16$ kHz sampling rate containing a single English word each [31]. The dataset contains words spoken by $1\thinspace{}864$ speakers. Specifically, we use the version $2$ of the dataset, consisting of $105\thinspace{}829$ audio files representing 35 classes. For dataset partitioning into training, testing, and validation sets, we employed the same hashing function used in [24].
@@
We also encode both the HD and SC datasets with LCA, selecting $k=700$ filters to match the number of channels in SHD and SSC. This allows us to compare, in Sec. IV, our work with state-of-the-art methods that use SHD/SSC with all the channels. We choose the values for the filter length $F_{l}$ and the stride $r$ that ensure to have similar analysis window (10 ms) as temporal histograms of SHD and SSC. We set the threshold $\lambda=0.00045$ for HD and $\lambda=0.0007$ for SC to maintain similar sparsity as SHD and SSC. The values used for the dictionary hyperparameters are in Tab. I. The representations of HD and SC with this algorithm are called LCA-HD and LCA-SC, respectively. From these LCA representations we derive ALCA representation by adapting the Gammachirps therefore creating two additional representations ALCA-HD and ALCA-SC. All representations are normalized to a 010-1 range, ensuring all values fall between 0 and $1$.
@@ III-A Static Features
Figure 3: Best accuracy achieved by fully connected ANNs with both HD and SC datasets with three representations: LAUSCHER S[HD||SC], LCA -[HD||SC], and ALCA-[HD||SC].
@@ III-B Importance of Temporal Features
where $_{\text{c}}(.)$ is the average number of non-zero coefficients in the considered representation (ALCA or LCA) computed from the test set of the considered dataset, and $_{\text{s}}(.)$ is the average spike count of the LAUSCHER representation computed from the test set of the considered dataset.
@@ III-D Spiking Activity
Quantized Model
Test Accuracy (%)
$83.77\pm 1.2$ $\boldsymbol{94.63\pm 0.8}$ $94.38\pm 1.0$ $68.53\pm 0.4$ $84.44\pm 0.5$ $\boldsymbol{86.40\pm 0.4}$
@@ III-E Power/Energy Efficiency on Loihi 2
The inference on the GPU utilizes (Lava-DL SLAYER), and its efficiency is measured using nvidia-smi on Linux with kernel version 5.4.0-7634, Python 3.8.1 and PyTorch 2.0.0. Loihi’s efficiency is measured using the power probes of Lava 0.9.0 on Oheogultch board ncl-ext-og-05.
@@
LAUSCHER [41] Feed-forward SNN with adaptive axonal delays $92.45$ -
LAUSCHER [42] Feed-forward SNN with spatio-temporal filters $92.36$ -
@@
LAUSCHER [43] RSNN with temporal attention $91.08$ -
LAUSCHER [44] Feed-forward SNN with random dendritic delays $90.88$ -